Circular No. 7727

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001ep IN NGC 1699 D. Hutchings and W. D. Li, University of California at Berkeley, report

the discovery by LOTOSS (cf. IAUC 7514) of an apparent supernova in unfiltered images taken with the Katzman Automatic Imaging Telescope (KAIT) on Oct. 3.5 (mag ∼ 17.1) and 4.4 UT (mag ∼ 16.8). SN 2001ep is located at α = 4^{h}57^{m}00^{s}.26, δ = −4^{o}45′40′′.2 (equinox 2000.0), which is 10′′.2 east and 15′′.8 south of the nucleus of NGC 1699. A KAIT image taken on Sept. 28.4 showed nothing at this position (limiting mag ∼ 19.0).

POSSIBLE NOVA IN CENTAURUS P. R. Holvorcem, Campinas, Brazil, provides the following precise po-

sition (measured using USNO-A2.0 reference stars) for the object reported on IAUC 7726, from his CCD images taken with a 0.36-m f/3.6 Schmidt-Cassegrain telescope in twilight and with the star at low altitude on Oct. 3.907 UT, when he found red mag 8.5: α = 13^{h}55^{m}41^{s}.27, δ = −64^{o}15′57′′.9 (equinox 2000.0). Holvorcem adds that this object does not appear on the Digitized Sky Survey. G. J. Garradd, Loomberah, New South Wales, reports the following position end figures (mean of three CCD images using 320 UCAC stars; total rms residuals 0′′.6) obtained with a 0.45-m f/5.4 reflector on Oct. 4.412: 41^{s}.23, 57′′.1. A very red star appears in the USNO-A2.0 catalogue with position end figures 41^{s}.25, 57′′.6 (red mag 17.6, blue mag 20.8). Visual magnitude estimates: Oct. 3.965 UT, 9.7 (A. Amorim, Florianopolis, Brazil); 4.540, 9.7 (A. Pearce, Nedlands, W. Australia).

V4740 SAGITTARII Visual magnitude estimates: Sept. 17.956 UT, 7.9 (J. G. de S. Aguiar,

Campinas, Brazil); 19.064, 8.0 (R. Shida, São Paulo, Brazil); 20.762, 8.3 (K. Hornoch, Lelekovice, Czech Republic); 25.584, 8.9 (Pearce); Oct. 3.975, 9.3 (Amorim).

COMET C/2000 WM_{1} (LINEAR) Total visual magnitude estimates: Aug. 16.99 UT, 13.8 (R. J. Bouma,

Groningen, The Netherlands, 0.31-m reflector); 25.40, 14.3 (A. Hale, Cloudcroft, NM, 0.41-m reflector); Sept. 15.99, 12.3 (J. Carvajal, Avila, Spain, 0.45-m reflector); 22.02, 12.4 (K. Sárneczky, Ágasvŕ, Hungary, 0.44-m reflector); 26.98, 11.5 (G. Sostero, Remanzacco, Italy, 0.45-m reflector); 29.13, 11.5 (Bouma, 0.25-m reflector).

2001 October 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green